OpenStudy (saifoo.khan):

How about this one? x4 – x3 + 7x2 – 9x – 18

OpenStudy (anonymous):

This one is actually easier

OpenStudy (saifoo.khan):

H.O.W.?

OpenStudy (anonymous):

substitute x by (u - B/(4A)), which is u - 1/4, and then expand

OpenStudy (saifoo.khan):

o_O No idea. LOL.

OpenStudy (anonymous):

expansion leads to \[u^4-2u^3+\frac{65u^2}{8}-\frac{51u}{4}-\frac{3915}{256}\]

OpenStudy (saifoo.khan):

lol, still no idea.

OpenStudy (anonymous):

\[x^4-x^3+7 x^2-9 x-18=(x-2) (x+1) \left(x^2+9\right) \]

OpenStudy (saifoo.khan):

H.O.W.?

OpenStudy (anonymous):

I did a mistake somewhere, since u^3 remains :-P

OpenStudy (anonymous):

I guess the method is complex :-P try doing the posible rational roots and trial & error method

OpenStudy (anonymous):

possible rational roots +-(1, 2, 3, 6, 9, 18)

OpenStudy (turingtest):

I would try to plug in low numbers and look for a zero. I would hopefully have tried 2 OR -1 which would have given me a zero and then divide the polynomial out

OpenStudy (anonymous):

why do you have these questions? your teacher wants you to burn :-D ?

OpenStudy (saifoo.khan):

Trial and error sounds the best. haha @adg

OpenStudy (anonymous):

im full now :p

OpenStudy (anonymous):

and the food tasted great! goodluck with ur problems seafood...

OpenStudy (saifoo.khan):

lol, thanks.

OpenStudy (saifoo.khan):

but when x = 1, im getting 20

OpenStudy (saifoo.khan):

-20

OpenStudy (anonymous):

From prior experience it appears that Mathematica can factor any expression that is factorable. I have no idea what factoring algorithms have been incorporated.

OpenStudy (saifoo.khan):

but when x = 1, im getting -20

OpenStudy (turingtest):

you tried x=-1 ? that's the root supposedly

OpenStudy (saifoo.khan):

damn!!

OpenStudy (saifoo.khan):

my bad. LOL

OpenStudy (saifoo.khan):

Thanks guys.

OpenStudy (turingtest):

you know how to divide polynomials right?

OpenStudy (saifoo.khan):

Yes i do know.

OpenStudy (saifoo.khan):

Both synthetic and another one.

OpenStudy (saifoo.khan):

Long division.

OpenStudy (turingtest):

then that's all I got

OpenStudy (saifoo.khan):

Wait, how can i solve this by that way? :D

OpenStudy (turingtest):

When you divide the polynomial by (x-1) you should get another cubic. Then you have to find that 2 is also a root. You could use Descartes' Rule of signs to tell you how many negative or positive roots a polynomial has, so using that could save you some time there. Once you divide out the (x-2) you have a nice quadratic we all love so much.